OpenStudy (anonymous):

|2z-7|=13 HELP absolute value is stressful

OpenStudy (anonymous):

Since absolute value makes anything positive we know that if 2z-7 = -13, the absolute value would make it positive 13. If 2z-7 = 13 then the absolute value would keep the answer same, that is to say it will remain to be 13. So here we have to cases, either 2z-7 = -13 or 2z - 7 = 13 but in either case the absolute value makes the answer positive which is 13. So now we solve for z in both cases to see the possible values that z could represent. 2z - 7 = -13 2z = -6 z = -3 OR 2z - 7 = 13 2z = 20 z = 10 Therefore, z = 10 or z = -3.

OpenStudy (anonymous):

Thinks learned alot your a LIFE SAVOR!!!!! :)

OpenStudy (anonymous):

Mind giving me a medal or becoming a fan? -.-